Find the Thevenin equivalent circuit at terminals a-b using simplified mesh analysis.و 15 V 120 بم 1A 4A (1)
Answer #1

L 1212 AA . 10.1 E 2-070 Rim - 2010 - 20 6678 20tro 12V 1 H otov Ven - 8 I - 12 I d5-do I-40=o -30I = 63 I=-2.1 Amp The 40+ [